OpenStudy (anonymous):

find all intergers x and y such that x^4-y^2=15

OpenStudy (anonymous):

\[(x^2+y)(x^2-y)=15\]

OpenStudy (experimentx):

woops!! didn't realize ^4

OpenStudy (anonymous):

\[x^2+y=5\] \[x^2-y=3\] ....or \[x^2+y=15\] \[x^2-y=1\]

OpenStudy (anonymous):

is that logical 15=3*5=15*1

OpenStudy (experimentx):

the factors of 15 are 1x15 and 3x5 -------------------------- for x^2 and y, it would yield 4 possible solutions. have to check system for integral values.

OpenStudy (experimentx):

you can do it other way x^2 + y = 3 and x^2-y=5 and same for other. just have to check for integral values.

OpenStudy (experimentx):

x^2 + y = 3 and x^2-y=5 x^2 + y = 3 x^2 - y = 5 ------------ 2x^2 = 8 x = +- 2 ------------ y = -1

OpenStudy (anonymous):

\[x^2=4,y=1\] \[x^2=16,y=1\]

OpenStudy (experimentx):

other way around x^2 + y = 5 x^2 - y = 3 ----------- x = +-2 and y=1

OpenStudy (anonymous):

thanks its working well

OpenStudy (experimentx):

x^2 + y = 1 x^2 - y = 15 ------------- x = +- 4, y = -15 x^2 + y = 15 x^2 - y = 1 ------------- x = +- 4, y = -1

OpenStudy (experimentx):

altogether 8 solutions!! since it was quadratic.

OpenStudy (raden):

@experimentX , if x^2+y = 1 and x^2-y = 15 ---------- + 2x^2 = 16 x^2 = 8 x=+-sqrt(8), not satisfying (because x must be an integer number) and for : x^2+y=15 x^2-y= 1 -------- + 2x^2 = 16 x^2 = 8 x=+-sqrt(8) not satisfying (because x must be an integer number) so, for x^2+y=1 and x^2-y=15 or reverse, no one satisfy for x integer number. so, only 4 solutions.

OpenStudy (experimentx):

woops!! sorry ... you are correct. I wasn't careful.